# Hybrid User Interfaces: Past, Present, and Future of Complementary Cross-Device Interaction in Mixed Reality

### v1

@@ 6.3 Evaluation, Assessments, and Toolkits
Brudy et al. [16] described that research on cross-device interaction can be considered as a “constructive problem” [106]. This is reflected in our corpus: The majority of papers focus on creating new artifacts. These papers “re-envision and push the boundaries of interaction possibilities” [16]—a common theme for research in HCI that is often “much better at proposing new technologies than at validating them” [64], especially when novelty is expected as a key contribution [^6] https://sigbed.org/2022/08/22/the-toxic-culture-of-rejection-in-computer-science/ last accessed on 2024-09-06. This further fragments research on HUIs, as artifact and empirical contributions tend to drift apart [16]. While some research addressed the call of Brudy et al. [16] for a frame of reference to compare cross-device interaction techniques [156], this is still missing for research on MR-HUIs. One possibility to create such frames of reference and further systematically study (and later compare) research within this space is to consider experiments as design-informing activities [107]. Here, different design alternatives (e.g., meaningful combinations of input and output components) can serve as independent variables for different use cases. This allows us to study their effect on general (e.g.,  time and error) and use case-specific dependent variables, such as utilization of devices [129]—focusing on the effect, influence, and utility of each combination, beyond a technical perspective [16].
@@
HUIs differ from single-device user interfaces by involving multiple interaction components, making assessments more complex. For example, cognitive workload (a typical metric in HCI user studies [84]) can be measured post-hoc using subjective questionnaires like NASA TLX [60, 59], but real-time, objective methods such as eye tracking are preferable [84]. As MR-HUIs often include an HWD, continuous assessment of eye movements via built-in sensors is achievable. However, switching between interaction components complicates data collection, as some interaction components might not support eye tracking or may have varying data quality, requiring data fusion or repeated calibration. This issue becomes more complex with multiple users frequently switching between interaction components. Although subjective questionnaires simplify user studies, they reduce data accuracy, as participants might forget specific details, especially in highly dynamic MR-HUIs. Objective, real-time assessments are critical for understanding MR-HUIs, but careful study designs are necessary to manage complexity in conducting and analyzing them.
@@
To address this complexity, research has suggested a variety of toolkits that can support this process for MR user studies [66, 116, 20, 103]. Yet, these toolkits also hold the potential to support the design, conduction, and analysis procedures of user studies involving MR-HUIs.
@@ 6.4 Authoring Mixed Reality Hybrid User Interfaces
Creating MR-HUIs can be challenging: On the one hand, designers require evidence-based guidelines that focus on the integration of multiple devices [30], such as determining the optimal ergonomic content distribution [157, 42]. On the other hand, developers need to deal with implementations on different platforms. Although web technologies (e.g., WebXR) can standardize development, their support on commercial HWDs is uncertain [17]. In addition, toolkits (e.g., for integration of multiple devices [51], interaction component configuration [123], and synchronization [65]) can help focus on the implementation at hand, rather than worrying about implementation details. Here, a common grammar (cf. VEGA [128, 127]) can be useful, allowing content to responsively adapt to each device automatically.
@@ 6.5 Exploring Holistic Real Life Applicability
The increase in artifacts and evaluations within our corpus indicates that the nascent space represented by our corpus is slowly maturing. Research prototypes and their interaction techniques have so far been investigated in isolation, yet “it is clear that [these] facets should not be discussed in isolation; instead, they are highly interconnected and affect each other” [86]. Given the wide applicability of existing (see Sec. 5.2) and future systems within our HUI subset (e.g., robotics [95], medical domain [115, 86, 135], explainable artificial intelligence [102]), we need to consider their role within the holistic context of their work environments and unveil their unique challenges and opportunities. Furthermore, consolidating common interaction techniques into a library of design patterns can help researchers and practitioners alike in the design and implementation of holistic MR-HUIs.
@@ 6.6 Collaboration
The majority of the papers in our corpus focus on single-user systems. Yet, some papers indicated a collaborative setting, either with a shared interaction component (e.g., a shared output) or using individual interaction components.
Regarding co-located collaboration, the choice whether there is a shared interaction component or individual interaction components can directly influence the type of collaboration: For example, Butscher et al. [18] used a large interactive tabletop display as a shared input device, combined with an AR HWD for each user. Limited by the touch input of the shared device, only one user is able to manipulate the content, thus enforcing a closely coupled collaboration [58, 142]. In contrast, providing users with individual interaction components enables loosely coupled, parallel activities [130]. Similarly, asymmetric device setups might benefit from individual device affordances [49] but also lead to an asymmetry of roles [155].
Lastly, no paper in our corpus features remote collaboration. In such cases, techniques known from MR remote collaboration could be applied: Virtual avatars can represent the remote person to create awareness [57, 56] while methods to align workspaces can allow for deictic referencing [48]. The use of diverse interaction components might require user representations per interaction component, and highly dynamic environments might benefit from transitional user representations [155].
@@ 7 The Future of Hybrid User Interfaces
So, what is a HUI now?  The term “hybrid” in HUI introduces ambiguity, as it could describe any combination of multiple interfaces and concepts. This ambiguity and divergent facilitation of the term is further demonstrated by contrasting the original definition of HUIs—referring to “heterogeneous display and interaction” [47] or “different interface” [43] technologies—with its most prevalent use today (i.e., combining 2D and MR components). This led to our decision to not sharply define HUIs and MR-HUIs because of the (1) historical developments and usages of the term, its (2) relations and overlaps to adjacent research areas, (3) the abundance of possibilities for combinations as shown in our taxonomy, and (4) potential edge cases. However, we opted to modernize Feiner and Shamash’s definition of HUIs [47] (see Sec. 2.3) while specifically focusing on the initial device combinations of 2D and MR interaction components.
@@
Through our work on this survey and our own experience, we can say that the term HUI (and with that MR-HUI) is more a kin to a fuzzy concept—“a collection of objects [that do] not have sharp, clear-cut boundaries” [21]. Thereby, HUIs can be characterized with certain attributes as described in Sec. 2.4. In the end, actual manifestations may vary—depending on the discussed ever-shifting window of opportunities—where HUI boundaries remain adaptable to emerging interaction paradigms, technological affordances, and contextual demands. However, if the term is adaptable in such a way, we must ask:
@@
Will there be HUIs in the future?  On the bright side, HUIs combine complementary device technologies, taking advantage of the strengths of each technology, creating user interfaces where “the real power [ …] comes not from any one of these devices; it emerges from the interaction of all of them” [149]. This combination can create flexible and modular systems in which the user can freely decide when to employ which type of feature (i.e., interaction component).
@@
However, from a rather critical perspective, HUIs could be considered as a compensation mechanism for the deficiencies of individual interaction components, such as the limited precision of gesture recognition in AR environments. By including additional technologies to address the shortcomings of any one input method, these systems can become too complex to design, evaluate, and eventually use. With advances in technology, the need for HUIs composed of multiple interaction components may diminish: An ultimate interface might not require the combination of multiple interaction components to compensate for individual weaknesses. Instead, at some point in the future, there will be self-contained “devices” that streamline interaction into a single cohesive system. This leads to the question:
@@
Is there an expiration date for HUIs?  We believe so but cannot say when. Over 30 years of research on HUIs and related research streams have shown a maturing of concepts and usage of technology. Although some of the pioneering works are often technology-driven, proving the feasibility of hardware-heavy combinations, newer works are more mindful of the usage of technology. At the same time, we observed an increase in various evaluation strategies over time (see Fig. 7). Both can be attributed to advances in technology (e.g., advanced built-in tracking capabilities)—minimizing the need for additional technical augmentation while focusing more on the user of such systems. This shifts the focus from purely technical work towards studying the individual benefits of various design alternatives (e.g., different variants of HUIs) and emphasizes the value and utility of individual interaction components and the user experience of HUIs. This maturing process led to current off-the-shelf devices (e.g., Apple Vision Pro [^7] https://apple.com/apple-vision-pro/, last accessed on 2024-09-06) and applications (e.g., Immersed [^8] https://immersed.com, last accessed on 2024-09-06), which already bring HUIs to the consumer market. With the maturation of concepts, an increase in empirical contributions, the advent of consumer-oriented HUIs, and the excelling development of other research areas such as artificial intelligence or even brain-computer interfaces, we can now take a step back and take a look at the bigger picture with our last question:
@@
What does the future of HUIs look like?  With our attributes, we identified a classification of past and current MR-HUIs. The interplay of interaction components is fundamental for HUIs: A typical HUI consists of multiple interaction components in heterogeneous roles, complementing each other, and forming a single application. Past and present HUIs often consider “interaction component” as a synonym for “device”—creating a cross-device interaction. However, by reconsidering what constitutes an interaction component, we can draw inspiration from concepts such as complementary interfaces [154] and related ideas [39]: These concepts are device-agnostic, opening up the research space for combinations of interaction components beyond devices and avoiding being trapped in technological restrictions. Here, new meaningful combinations of interaction components that include different input and output modalities, combinations of implicit and explicit interaction, and various interaction techniques provide an avenue of research opportunities. Thus, we can expect a change in how HUIs will be designed, built, and evaluated until they eventually disappear when approaching the “ultimate display” [140].
@@ Acknowledgements
We thank all participants of the IEEE ISMAR 2023 Workshop on Hybrid User Interfaces.
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy: EXC-2068, 390729961 – Cluster of Excellence “Physics of Life” and EXC 2050/1, 390696704 – Cluster of Excellence “Centre for Tactile Internet” (CeTI) of TU Dresden; by DFG grant 389792660 as part of TRR 248 – CPEC; and by Project-ID 251654672 – TRR 161. This work was supported by the German Federal Ministry of Education and Research (BMBF, SCADS22B) and the Saxon State Ministry for Science, Culture and Tourism (SMWK) by funding the competence center for Big Data and AI ”ScaDS.AI Dresden/Leipzig”; by the Alexander von Humboldt Foundation (BMBF); by the Independent Research Fund Denmark (DFF) (grant ID: 10.46540/3104-00008B); and by Villum Investigator grant VL-54492 by Villum Fonden.

### v2

@@ 6.3 Evaluation, Assessments, and Toolkits
Brudy et al. [brudy2019crossdevice] described that research on cross-device interaction can be considered as a “constructive problem” [oulasvirta2016hci]. This is reflected in our corpus: The majority of papers focus on creating new artifacts. These papers “push the boundaries of interaction possibilities” [brudy2019crossdevice]—a common theme for research in HCI that is often “much better at proposing new technologies than at validating them” [hornbaek2011whys], especially when novelty is expected as a key contribution [^6] https://sigbed.org/2022/08/22/the-toxic-culture-of-rejection-in-computer-science/ last accessed on 2025-12-01. This further fragments research on HUIs, as artifact and empirical contributions tend to drift apart [brudy2019crossdevice]. While some research addressed the call for a frame of reference [brudy2019crossdevice] to compare cross-device interaction techniques [zagermann2020otherhand], this is still missing for research on MR-HUIs. One possibility to create such frames of reference and systematically study (and later compare) research within this space is to consider experiments as design-informing activities [oulasvirta2022counterfactual]. Here, different design alternatives (e.g., meaningful combinations of input and output components) can serve as independent variables for different use cases. This allows us to study their effect on general (e.g.,  time, error) and use case-specific dependent variables, such as utilization of devices [schroder2023model]—focusing on the effect, influence, and utility of each combination, beyond a technical perspective [brudy2019crossdevice].
@@
HUIs differ from single-device user interfaces by involving multiple IO components, making assessments more complex. For example, cognitive workload (a typical metric in HCI user studies [kosch2023survey]) can be measured post-hoc using subjective questionnaires like NASA TLX [hart1988development, hart2006nasatask], but real-time, objective methods such as eye tracking are preferable [kosch2023survey]. As MR-HUIs often include an HWD, continuous assessment of eye movements via built-in sensors is achievable. However, switching between IO components complicates data collection, as some IO components might not support eye tracking or may have varying data quality, requiring data fusion or repeated calibration. This issue becomes more complex with multiple users frequently switching between IO components. Although subjective questionnaires simplify user studies, they reduce data accuracy, as participants might forget specific details, especially in highly dynamic MR-HUIs. Objective, real-time assessments are critical for understanding MR-HUIs, but careful study designs are necessary to manage complexity in conducting and analyzing them.
@@
To address this complexity, research has suggested a variety of toolkits that can support this process for MR user studies [hubenschmid2022relive, reipschlager2022avatar, buschel2021miria, nebeling2020mrat]. Yet, these toolkits also hold the potential to support the design, conduction, and analysis procedures of user studies involving MR-HUIs.
@@ 6.4 Authoring Mixed Reality Hybrid User Interfaces
Creating MR-HUIs can be challenging: Designers require evidence-based guidelines that focus on the integration of multiple devices [daeijavad2023hybrid], such as determining the optimal ergonomic content distribution [zaky2023opportunities, belo2021xrgonomics]. Yet, developers also need to deal with implementations on different platforms. Although web technologies (e.g., WebXR) can standardize development, their support on commercial HWDs is uncertain [butcher2023don]. In addition, toolkits (e.g., for integration of multiple devices [gonzalez2024xdtk], IO component configuration [sandor2005immersive], synchronization [hubenschmid2023colibri]) can help focus on the implementation at hand, rather than worrying about implementation details. Here, a common grammar (cf. VEGA [satyanarayan2016reactive, satyanarayan2017vegalite]) could allow content to responsively adapt to each device.
@@ 6.5 Exploring Holistic Real Life Applicability
The increase in artifacts and evaluations within our corpus indicates that the nascent space represented by our corpus is slowly maturing. Research prototypes and their interaction techniques have so far been investigated in isolation, yet “it is clear that [these] facets should not be discussed in isolation; instead, they are highly interconnected and affect each other” [krug2023discussing]. Given the wide applicability of existing (see Sec. 5.2) and future systems within our HUI subset (e.g., robotics [lunding2024proposing], medical domain [reinschluessel2023exploring, krug2023discussing, smith2023integrating], explainable artificial intelligence [mendez2023how]), we need to consider their role within the holistic context of their work environments and unveil their unique challenges and opportunities. Furthermore, consolidating common interaction techniques into a library of design patterns can help researchers and practitioners alike in the design and implementation of holistic MR-HUIs.
@@ 6.6 Collaboration
The majority of the papers in our corpus focus on single-user systems. Yet, some papers indicated a collaborative setting, either with a shared IO component (e.g., a shared output) or using individual IO components.
Regarding co-located collaboration, the choice whether there is a shared IO component or individual IO components can directly influence the type of collaboration: For example, Butscher et al. [butscher2018clusters] used a large interactive tabletop display as a shared input device, combined with an AR HWD for each user. Limited by the touch input of the shared device, only one user is able to manipulate the content, thus enforcing a closely coupled collaboration [gutwin1998design, tang2006collaborative]. In contrast, providing users with individual IO components enables loosely coupled, parallel activities [schroder2023collaborating]. Similarly, asymmetric device setups might benefit from individual device affordances [friedl-knirsch2023exploring] but also lead to an asymmetry of roles [zagermann2023challenges].
Lastly, no papers in our corpus feature remote collaboration. In such cases, techniques known from MR remote collaboration could be applied: Virtual avatars can represent the remote person to create awareness [gronbaek2024blended, gronbaek2023transitional] while methods to align workspaces can allow for deictic referencing [fink2022relocations]. The use of diverse IO components might require user representations per IO component, and highly dynamic environments might benefit from transitional user representations [zagermann2023challenges].
@@ 7 The Future of Hybrid User Interfaces
So, what is a HUI now?  The term “hybrid” in HUI introduces ambiguity, as it could describe any combination of multiple interfaces and concepts. This ambiguity and divergent facilitation of the term is further demonstrated by contrasting the original definition of HUIs—referring to “heterogeneous display and interaction” [feiner1991hybrid] or “different interface” [feiner1994redefining] technologies—with its most prevalent use today (i.e., combining 2D and MR components). This led to our decision to not sharply define HUIs and MR-HUIs because of the (1) historical developments and usages of the term, its (2) relations and overlaps to adjacent research areas, (3) the abundance of possibilities for combinations as shown in our taxonomy, and (4) potential edge cases. However, we opted to modernize Feiner and Shamash’s definition of HUIs [feiner1991hybrid] (see Sec. 2.3) while specifically focusing on the initial device combinations of 2D and MR IO components.
@@
Through our work on this survey and our own experience, we can say that the term HUI (and with that MR-HUI) is more akin to a fuzzy concept—“a collection of objects [that do] not have sharp, clear-cut boundaries” [belohlavek2011concepts]. Thereby, HUIs can be characterized with certain attributes as described in Sec. 2.4. In the end, actual manifestations may vary—depending on the discussed ever-shifting window of opportunities—where HUI boundaries remain adaptable to emerging interaction paradigms, technological affordances, and contextual demands. However, if the term is adaptable in such a way, we must ask:
@@
Will there be HUIs in the future?  On the bright side, HUIs combine complementary device technologies, taking advantage of the strengths of each technology, creating user interfaces where “the real power [ …] comes not from any one of these devices; it emerges from the interaction of all of them” [weiser1999computer]. This combination can create flexible and modular systems in which the user can freely decide when to employ which type of feature (i.e., IO component).
@@
However, from a rather critical perspective, HUIs could be considered as a compensation mechanism for the deficiencies of individual IO components, such as the limited precision of gesture recognition in AR environments. By including additional technologies to address the shortcomings of any one input method, these systems can become too complex to design, evaluate, and eventually use. With advances in technology, the need for HUIs composed of multiple IO components may diminish: An ultimate interface might not require the combination of multiple IO components to compensate for individual weaknesses. Instead, at some point in the future, there will be self-contained “devices” that streamline interaction into a single cohesive system. This leads to the question:
@@
Is there an expiration date for HUIs?  We believe so, but cannot say when. Over 30 years of research on HUIs and related research streams have shown a maturing of concepts and usage of technology. Although some of the pioneering works are often technology-driven, proving the feasibility of hardware-heavy combinations, newer works are more mindful of the usage of technology. At the same time, we observed an increase in various evaluation strategies over time. Both can be attributed to advances in technology (e.g., advanced built-in tracking capabilities)—minimizing the need for additional technical augmentation while focusing more on the user of such systems. This shifts the focus from purely technical work towards studying the individual benefits of various design alternatives (e.g., different variants of HUIs) and emphasizes the value and utility of individual IO components and the user experience of HUIs. This maturing process led to current off-the-shelf devices (e.g., Apple Vision Pro [^7] https://apple.com/apple-vision-pro/, last accessed on 2025-12-01) and applications (e.g., Immersed [^8] https://immersed.com, last accessed on 2025-12-01), which already bring HUIs to the consumer market. With the maturation of concepts, an increase in empirical contributions, the advent of consumer-oriented HUIs, and the development of research areas such as artificial intelligence or even brain-computer interfaces, we can take a step back and look at the bigger picture with a last question:
@@
What does the future of HUIs look like?  With our attributes, we identified a classification of past and current MR-HUIs. The interplay of IO components is fundamental for HUIs: A typical HUI consists of multiple IO components in heterogeneous roles, complementing each other, and forming a single application. Past and present HUIs often consider “IO component” as a synonym for “device”—creating a cross-device interaction. However, by reconsidering what constitutes an IO component, we can draw inspiration from concepts such as complementary interfaces [zagermann2022complementary] and related ideas [elmqvist2023anywhere]: These concepts are device-agnostic, opening up the research space for combinations of IO components beyond devices and avoiding being trapped in technological restrictions. Here, new meaningful combinations of IO components that include different input and output modalities, combinations of implicit and explicit interaction, and various interaction techniques provide an avenue of research opportunities. Thus, we can expect a change in how HUIs will be designed, built, and evaluated until they eventually disappear when approaching the “ultimate display” [sutherland1965ultimate].
@@ Acknowledgements
We thank all participants of the IEEE ISMAR 2023 Workshop on Hybrid User Interfaces. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy: EXC 2050/1, 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TU Dresden; by DFG grant 389792660 as part of TRR 248 – CPEC; and by Project-ID 251654672 – TRR 161. This work was supported by the German Federal Ministry of Education and Research (BMBF, SCADS22B) and the Saxon State Ministry for Science, Culture and Tourism (SMWK) by funding the competence center for Big Data and AI ”ScaDS.AI Dresden/Leipzig”; by the Alexander von Humboldt Foundation (BMBF); by the Independent Research Fund Denmark (DFF) (grant ID: 10.46540/3104-00008B); and by Villum Investigator grant VL-54492 by Villum Fonden.